OpenStudy (anonymous):

If sin x=1/2 , what is cos x?

OpenStudy (anonymous):

one possibility is \(\frac{\sqrt{3}}{2}\) but it could be negative

OpenStudy (anonymous):

look at the picture of the unit circle on the last page of the cheat sheet, and see where the second coordinate is \(\frac{1}{2}\) then look at the first coordinate